# Dataset Distillation as Pushforward Optimal Quantization

### v2

@@ Abstract
Dataset distillation aims to find a synthetic training set such that training on the synthetic data achieves similar performance to training on real data, with orders of magnitude less computational requirements. Existing methods can be broadly categorized as either bi-level optimization problems that have neural network training heuristics as the lower level problem, or disentangled methods that bypass the bi-level optimization by matching distributions of data. The latter method has the major advantages of speed and scalability in terms of size of both training and distilled datasets. We demonstrate that when equipped with an encoder-decoder structure, the empirically successful disentangled methods can be reformulated as an optimal quantization problem, where a finite set of points is found to approximate the underlying probability measure by minimizing the expected projection distance. In particular, we link existing disentangled dataset distillation methods to the classical optimal quantization and Wasserstein barycenter problems, demonstrating consistency of distilled datasets for diffusion-based generative priors. We propose Dataset Distillation by Optimal Quantization, based on clustering in a latent space. Compared to the previous SOTA method D4M, we achieve better performance and inter-model generalization on the ImageNet-1K dataset with trivial additional computation, and SOTA performance in higher image-per-class settings. Using the distilled noise initializations in a stronger diffusion transformer model, we obtain SOTA distillation performance on ImageNet-1K and its subsets, outperforming diffusion guidance methods.
@@ 1 Introduction
Training powerful neural networks requires a large amount of data, and thus induces high computational requirements. Dataset distillation (DD) targets this computational difficulty by changing the data, as opposed to other parts of training such as optimization or architecture [53]. The DD objective consists of finding a synthetic training set, such that training a neural network on the synthetic data yields similar performance.
There are several closely related notions of reducing computational load when training new models on datasets. Core-set methods find a subset of training data (as opposed to synthetic data) such that a model trained on said subset will have similar performance [30, 11]. Model distillation, sometimes known as knowledge distillation, aims to train a smaller (student) model that predicts the output of a pre-trained (teacher) model [12, 40]. Importance sampling methods accelerate training by weighting training data, finding examples that are more influential for training [36]. For more detailed surveys on dataset distillation methods and techniques, we refer to [57, 44].
@@ 1.1 Bi-level formulation of dataset distillation
We first present the DD problem statement as given in [44]. Denote a training set by $\mathcal{T}$ (more generally, distribution of training data), and the expected and empirical risks (test and training loss) by $\mathcal{R}$ and $\mathcal{L}$ respectively, evaluated for a network parameter $\Phi$. Given a learning algorithm $\Phi=\Phi_{S}$, which takes a dataset $S$ and returns an $\mathcal{L}$-optimal parameter, the goal of DD is to find a synthetic dataset $\mathcal{S}$ (of given size) minimizing the test loss discrepancy:
$\mathcal{S}=\operatorname*{arg\,min}_{S}\left|\mathcal{R}(\Phi_{S})-\mathcal{R}(\Phi_{\mathcal{T}})\right|.$ (1)
@@
This formulation is computationally intractable. Approximations include replacing the minimum discrepancy objective with maximum test performance, replacing the learning algorithm $\Phi$ with an inner neural network optimization problem, and solving the outer minimization problem using gradient methods. Common heuristic relaxations to the bi-level formulation 1 include meta-learning [53, 10], distribution matching [58], and trajectory matching [5]. Other methods include neural feature matching [59, 28] and the corresponding neural tangent kernel methods [33, 32], representative matching [26], and group robustness [52]. For better scaling, [5, 31] consider using generative priors such as GANs to generate more visually coherent images, replacing the need for optimization with a neural network inversion task and increasing performance.
While the bi-level formulation follows naturally from the qualitative problem statement of dataset distillation, there are two main drawbacks, namely computational complexity and model architecture dependence. This limits the applicability to large-scale datasets and imposes additional requirements when using the distilled data.
The first prototypical example of using a generative model to alleviate the aforementioned drawbacks of the bi-level formulation is by “factorizing” the high-dimensional image-space optimization. This consists of jointly training generative “hallucinator” networks to generate samples from a common set of learnable “base” points [25]. This was able to increase cross-architecture performance when compared to previous image-space methods.
1.2 Disentangled methods
@@
The dimensionality of the underlying bi-level optimization problems limits the applicability on large scale datasets, which are particularly useful for computationally limited applications. For example, the ImageNet-1K dataset consists of 1.2M training images, totalling over 120GB of memory [9]. The full dataset ImageNet-21K consists of over 14M images and takes up around 1.2TB of memory, which is generally infeasible to train expert models on, and makes backpropagation through network training steps impossible.
@@
[56] is the first work to “disentangle” the bi-level optimization framework into three separate problems, named Squeeze, Recover and Relabel (SRe2L). In particular, the inner neural network optimization problem is replaced with matching statistics of batch-normalization layers. Curriculum Data Augmentation (CDA) uses adaptive training to get more performance [55]. [24] considers optimizing images such that neural network features are close to Wasserstein barycenters of the training image features. [51] proposes Realistic Diverse and Efficient Dataset Distillation (RDED), which replaces the latent clustering objective with a patch-based adversarial objective. [50] considers clustering directly in the latent space of a latent diffusion model (LDM) [43], named Dataset Distillation via Disentangled Diffusion Model (D4M). This avoids backpropagation when distilling and has constant memory usage with respect to images per class (IPC), a direct advantage over the linear memory scaling of optimization-based methods. Recent state-of-the-art methods consider fine-tuning diffusion models to have better statistics [14], and guiding the diffusion to maximize the influence of the distilled points [6], similarly to active learning.
@@
While dataset distillation has had extensive experimental effort, few proper theoretical justifications or computable formalizations exist in the literature. [44] proposes a high-level formulation based on minimizing the difference in test loss between learning on the full dataset versus the synthetic dataset. [22] considers dataset distillation as dependent on the desired inference task (typically classification with cross-entropy loss for image data). This allows them to interpret trajectory-matching as a mean-field control problem, and find analogous procedures for non-imaging examples like finding collocation points for training physics informed neural networks. However, they do not consider the problem of consistency or convergence of the distilled datasets, which we aim to address in this work.
We interpret dataset distillation as the task of finding a surrogate dataset that minimizes the Wasserstein distance with the original training dataset. As a consequence for training neural networks, the gradient steps for a well-distilled dataset will be similar to the steps for the training dataset. This generalizes the objectives of the bi-level approaches to allow for different network architectures.
1.3 Contributions
@@
    Disentangled dataset distillation can be theoretically justified using classical notions of optimal quantization and Wasserstein distance. Motivated by the empirical usage of clustering in latent spaces, we show in Theorem˜1 that optimal quantizations induce convergent approximations of gradients of population risk, and furthermore the rate is given by $\mathcal{O}(K^{-1/d})$, where $d$ is the dimension of the latent space and $K$ is the number of quantization points. This motivates the usage of a low-dimensional latent space.
@@
    From the aforementioned classical literature, we propose Dataset Distillation by Optimal Quantization (DDOQ) in Algorithm˜2, based on clustering in a low-dimensional latent space. This is related to a recent disentangled method D4M. Our proposed method has a smaller Wasserstein distance between the distilled latent points and latent data distribution. This theoretically results in network gradients that more closely approximate those from training on the full training dataset.
@@
    We algorithmically compare our proposed method with D4M and various baselines on the ImageNet-1K dataset using the same generative diffusion model backbone, demonstrating significantly better classification accuracy at varied IPC budgets and better cross-architecture generalization. Moreover, using the stronger diffusion transformer backbone, we provide a central comparison with SOTA disentangled and diffusion-based distillation methods, yielding competitive or better results than existing state-of-the-art on ImageNet-1K and some subsets.
This paper is structured as follows. Section˜2 covers related background and convergence results, including optimal quantization and diffusion models. Theorem˜1 in Section˜3 demonstrates consistency of the optimal quantizers when passed through diffusion-based generative priors to the image space, and proposes a simple extension of the D4M method by adding automatically-learned weights in the prototyping phase. We further fully explain the data distillation pipeline in a sequential and systematic manner from distillation, to training new models using the computed weights. Section˜4 contains experiments of our proposed method against the previously SOTA D4M method as well as other recent SOTA baselines on the large-scale ImageNet-1K dataset.
@@ 2 Background
This section covers the necessary background and concepts required for the theoretical interpretation of DD. Section˜2.1 first covers necessary notation. Section˜2.2 introduces the optimal quantization problem, the concept of quadratic distortion, and links it to Wasserstein optimality. Section˜2.3 provides a short exposition of score-based diffusion models, covering well-posedness and classical divergence bounds.
2.1 Definitions and Notation
Define $\mathcal{P}_{2}(\mathbb{R}^{d})$ to be the set of probability measures on $\mathbb{R}^{d}$ with finite second moment, not necessarily admitting a density with respect to the Lebesgue measure. We use ${\mathcal{W}}_{2}$ to denote the Wasserstein-2 distance between two probability distributions in $\mathcal{P}_{2}(\mathbb{R}^{d})$ [45], defined as
@@ 2.2 Optimal Quantization
For a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with finite second moments, an optimal quantization (or vector quantization) at level $K$ is a set of points $\{x_{1},...,x_{K}\}\subset\mathbb{R}^{d}$ such that the $\mu$-averaged Euclidean distance to the quantized points is minimal. This can be formulated as the minimizer of the (quadratic) distortion, defined as follows.
@@
Note that the assumption that $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ has finite second moments implies that the quadratic distortion is finite for any set of points, and that an optimal quantization exists [34]. We note that the optimal quantization weights are uniquely determined by the quantization points. This gives equivalence of the distortion minimization problem to the Wasserstein minimization problem, when restricted to measures of finite support [34].
@@
Suppose we have a quantization ${\mathbf{x}}=\{x_{1},...,x_{K}\}$. Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi clusters $\mu(\partial C_{i})=0$. Then the measure $\nu$ that minimizes the Wasserstein-2 distance 2 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$. Moreover, the optimal coupling is given by the projection onto the centroids.
For a given quantization $\mathbf{x}$} or quantization level $K$, the quadratic distortion satisfies respectively:
In other words, minimizing the quadratic distortion is equivalent to finding a Wasserstein-2 optimal approximation with a $K$-finitely supported (probability) measure.
@@
The case where the approximating measure is a uniform Dirac mixture is called the Wasserstein barycenter problem [8]. The Wasserstein error of the Wasserstein barycenter is higher than that of the optimal quantization, but it admits an easily computable dual representation.
Optimal quantizers converge in Wasserstein distance at a rate $\Theta(K^{-1/d})$ to their respective distributions as the number of particles $K$ increases [13]. Moreover, any limit point of the optimal quantizers is an optimal quantizer of the limiting distribution [41]. Moreover, it can be shown that if a sequence of distributions converges in Wasserstein distance $\mu_{n}\rightarrow\mu$, then so do the errors in quantization for a fixed quantization level [27, Thm. 4].
The quantizer can also be shown to have nice approximation properties when taking expectations of functions. A prototypical example for DD would have $f$ be the gradient of a neural network with respect to some loss function. This implies that a data distribution $\mu$ and its quantization $\nu$ induce similar training dynamics.
@@ 2.2.1 Solving the optimal quantization problem
To find an optimal quantizer, we now introduce a commonly used quantization algorithm: the competitive learning vector quantization (CLVQ) algorithm [1]. This is derived directly from gradient descent on the quadratic distortion, starting from an initialization in the convex hull of the support of $\mu$. For a given quantization, the gradient of the distortion has a representation in terms of $\mu$-centroids of the corresponding Voronoi cells, given explicitly in Section˜C.3.
The CLVQ algorithm is presented in Algorithm˜1. It consists of iterating: (i) sampling from $\mu$, (ii) computing the nearest cluster centroid, and (iii) updating the cluster centroid with weighted average. This algorithm is equivalent to the mini-batch $k$-means method when the stepsizes $\gamma_{i}$ are chosen to be the reciprocals of the number of points per-cluster [46, 37]. We can thus interpret mini-batch $k$-means as finding a local minima of the optimal quantization problem.
While this algorithm produces points ${\mathbf{x}}=(x_{1},...,x_{K})$, it remains to compute the associated weights approximating the measures of the Voronoi cells $\mu(C_{i})$ as in Proposition˜1. This can be done in an online manner within the same iterations [34].
1 Initialize weights ${\mathbf{w}}=(w_{1},...,w_{K})=(1/K,...,1/K)$;
2 while not converged do
3   Sample $X_{i}\sim\mu$;
7 $i\leftarrow i+1$;
8
9 end while
Result: quantization $\nu_{K}=\sum_{k=1}^{K}w_{i}\delta(x_{k}^{*})$
Algorithm 1 CLVQ
Proposition 3 ([3, Prop. 7]).
Assume that the measure $\mu\in{\mathcal{P}}_{2+\eta}(\mathbb{R}^{d})$ for some $\eta>0$, and that it assigns no mass to hyperplanes. Assume further that the grids ${\mathbf{x}}^{(t)}$ produced by CLVQ Algorithm˜1 converge to a stationary grid $\mathbf{x}$$^{*}$, i.e. $\nabla{\mathcal{G}}({\mathbf{x}}^{*})=0$, and that the step-sizes satisfy $\sum\gamma_{k}=+\infty$ and $\sum\gamma_{k}^{1+\delta}<\infty$ for some $\delta>0$. Then,

### v3

@@ Abstract
Dataset distillation aims to find a small synthetic training set, such that training on the synthetic data achieves similar performance to training on a larger training dataset. Early methods solve this by interpreting the distillation problem as a bi-level optimization problem. On the other hand, disentangled methods bypass pixel-space optimization by matching data distributions and using generative techniques, leading to better computational complexity in terms of size of both training and distilled datasets. We demonstrate that by using latent spaces, the empirically successful disentangled methods can be reformulated as an optimal quantization problem, where a finite set of points is found to approximate the underlying probability measure. In particular, we link disentangled dataset distillation methods to the classical problem of optimal quantization, and are the first to demonstrate consistency of distilled datasets for diffusion-based generative priors. We propose Dataset Distillation by Optimal Quantization (DDOQ), based on clustering in the latent space of latent diffusion models. Compared to a similar clustering method D4M, we achieve better performance and inter-model generalization on the ImageNet-1K dataset using the same model and with trivial additional computation, achieving SOTA performance in higher image-per-class settings. Using the distilled noise initializations in a stronger diffusion transformer model, we obtain competitive or SOTA distillation performance on ImageNet-1K and its subsets, outperforming recent diffusion guidance methods.
@@ 1 Introduction
Training powerful neural networks requires a large amount of data, and thus induces high computational requirements. Dataset distillation (DD) targets this computational difficulty by changing the data, as opposed to other parts of training such as optimization or architecture (Wang et al., 2018). The DD objective consists of finding a synthetic training set, such that training a neural network on the synthetic data yields similar performance.
There are several closely related notions of reducing computational load when training new models on datasets. Core-set methods find a subset of training data (as opposed to synthetic data) that achieve good training performance (Mirzasoleiman et al., 2020; Feldman, 2020). Model distillation, sometimes known as knowledge distillation, aims to train a smaller model that predicts the output of a larger model (Gou et al., 2021; Polino et al., 2018). Importance sampling methods accelerate training by weighting training data, finding examples that are more influential for training (Paul et al., 2021). For more detailed surveys on dataset distillation methods and techniques, we refer to (Yu et al., 2023; Sachdeva and McAuley, 2023).
Training data distributionLow-dimensional latent spaceDistilled latent points and weightsDistilled data distribution $\{(\quad,\quad)\}$(1)Encoding(2)Clustering(3)Decoding
Figure 1: Sketch of the proposed method pipeline. Using an encoder/decoder model, we map our high dimensional data to a low-dimensional space, which is then clustered using $k$-means. The clustered latent points and weights are then decoded to obtain the distilled data. This work argues that the weights are important when decoding; furthermore, “disentangled” distillation using an encoder-cluster-decoder framework is asymptotically consistent.
@@ 1.1 Bi-level formulation of dataset distillation
Denote a training set (more generally, distribution of training data) by $\mathcal{T}$, and the expected and empirical risks (test and training loss) by $\mathcal{R}$ and $\mathcal{L}$ respectively, evaluated for some parameter $\theta$. The goal of DD is to find a synthetic dataset $\mathcal{S}$ (of given size) minimizing the test loss discrepancy (Sachdeva and McAuley, 2023):
$\mathcal{S}=\operatorname*{arg\,min}_{S}\left|\mathcal{R}(\operatorname*{arg\,min}_{\theta}{\mathcal{L}}(S))-\mathcal{R}(\operatorname*{arg\,min}_{\theta}{\mathcal{L}}({\mathcal{T}}))\right|.$ (1)
@@
This formulation is computationally intractable. Approximations include replacing the minimum discrepancy objective with maximum test performance, replacing the learning algorithm $\Phi$ with an inner neural network optimization problem, and solving the outer minimization problem using gradient methods. Common heuristic relaxations to the bi-level formulation 1 include meta-learning (Wang et al., 2018; Deng and Russakovsky, 2022), distribution matching (Zhao and Bilen, 2023), and trajectory matching (Cazenavette et al., 2022). Other methods include neural feature matching (Zhou et al., 2022; Loo et al., 2022) and the corresponding neural tangent kernel methods (Nguyen et al., 2021; 2020), representative matching (Liu et al., 2023b), and group robustness (Vahidian et al., 2024). For better scaling, Cazenavette et al. (2022); Moser et al. (2024) consider using generative priors such as GANs to generate more visually coherent images, increasing performance and replacing the need for optimization with a neural network inversion task.
@@
While the bi-level formulation follows naturally from the qualitative problem statement of dataset distillation, there are two main drawbacks, namely computational complexity and model architecture dependence. The dimensionality of the underlying optimization problems limit the applicability on large scale datasets, which are particularly useful for computationally limited applications. For example, the ImageNet-1K dataset consists of 1.2M training images, totalling over 120GB of memory (Deng et al., 2009). The full dataset ImageNet-21K consists of over 14M images and takes up around 1.2TB of memory, which is generally infeasible to train expert models on, and makes backpropagation through network training steps impossible.
1.2 Disentangled and diffusion methods
@@
Yin et al. (2023) is the first work to “disentangle” the bi-level optimization framework into three separate problems, named Squeeze, Recover and Relabel (SRe2L). In particular, the inner neural network optimization problem is replaced with matching statistics of batch-normalization layers. Curriculum Data Augmentation (CDA) uses adaptive training to get more performance (Yin and Shen, 2024). Liu et al. (2023a) considers optimizing images such that neural network features are close to Wasserstein barycenters of the training image features. Sun et al. (2024) proposes Realistic Diverse and Efficient Dataset Distillation (RDED), which replaces the latent clustering objective with a patch-based adversarial objective. Su et al. (2024) considers clustering directly in the latent space of a latent diffusion model (LDM) (Rombach et al., 2022), named Dataset Distillation via Disentangled Diffusion Model (D4M). This avoids backpropagation when distilling and has constant memory usage with respect to images per class (IPC), a direct advantage over the linear memory scaling of optimization-based methods. Recent state-of-the-art methods consider fine-tuning diffusion models to have better statistics (Gu et al., 2024), and guiding the diffusion to maximize the influence of the distilled points (Chen et al., 2025), similarly to active learning.
@@
While dataset distillation has had extensive experimental effort, few theoretical justifications or computable formalizations exist in the literature. Sachdeva and McAuley (2023) proposes a high-level formulation based on minimizing the difference in test loss between learning on the full dataset versus the synthetic dataset. Kungurtsev et al. (2024) considers dataset distillation as dependent on the desired inference task (typically classification with cross-entropy loss for image data), interpreting trajectory matching as a mean-field control problem. No prior literature on dataset distillation addresses theoretically whether or not the distilled datasets are reasonable approximations of the input training data distribution. In this work, we address convergence in measure space for disentangled methods, demonstrating consistency and convergence of the distilled datasets.
@@
    We theoretically justify the disentangled dataset distillation framework, exploiting its structure to show that these methods converge to the true data distribution as the number of distilled points increases, using classical notions of optimal quantization and Wasserstein distance. Motivated by the empirical usage of clustering in latent spaces, we show in Theorem˜1 that optimal quantizations induce convergent approximations of gradients of population risk. Furthermore, Corollary˜1 shows the approximation rate is given by $\mathcal{O}(K^{-1/d})$, where $d$ is the dimension of the latent space and $K$ is the number of quantization points. This motivates the usage of a low-dimensional latent space to model the data distribution.
@@
    We propose Dataset Distillation by Optimal Quantization (DDOQ) in Algorithm˜1, a DD algorithm based on clustering in a low-dimensional latent space. Compared to a recent disentangled method D4M, our proposed method has a smaller Wasserstein distance between the distilled latent points and latent data distribution, indicating better approximation.
@@
    We algorithmically compare our proposed method with D4M and various common disentangled baselines on the ImageNet-1K dataset using the same generative diffusion model backbone, demonstrating significantly better classification accuracy at varied IPC budgets and better cross-architecture generalization. To demonstrate the potential of DDOQ, we additionally use the stronger diffusion transformer (DiT) backbone, used in recent diffusion-based methods. We provide a central comparison with SOTA disentangled and diffusion-based distillation methods, yielding competitive or better results than existing SOTA methods on ImageNet-1K and its subsets.
This work is structured as follows. Section˜2 covers related background and convergence results, including optimal quantization and diffusion models. Theorem˜1 in Section˜3 demonstrates consistency of the optimal quantizers when passed through diffusion-based generative priors to the image space, and motivates adding automatically-learned weights in the prototyping phase. Section˜3.1 details the data distillation pipeline in a sequential manner, from distillation to training new models using the computed weights. Section˜4 contains experiments of our proposed method against the previously SOTA D4M method as well as other recent SOTA baselines on the large-scale ImageNet-1K dataset.
@@ 2 Background
Define $\mathcal{P}_{2}(\mathbb{R}^{d})$ to be the set of probability measures on $\mathbb{R}^{d}$ with finite second moment, not necessarily admitting a density with respect to the Lebesgue measure. We use ${\mathcal{W}}_{2}$ to denote the Wasserstein-2 distance between two probability distributions in $\mathcal{P}_{2}(\mathbb{R}^{d})$ (Santambrogio, 2015).
@@ 2.1 Optimal Quantization
For a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, an optimal quantization (or vector quantization) at level $K$ is a set of points $\{x_{1},...,x_{K}\}\subset\mathbb{R}^{d}$ such that the $\mu$-averaged Euclidean distance to the quantized points is minimal. This can be formulated as the minimizer of the (quadratic) distortion, defined as follows.
@@
Note that the assumption that $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ has finite second moments implies that the quadratic distortion is finite for any set of points, and that an optimal quantization exists (Pagès, 2015). We note that the optimal quantization weights are uniquely determined by the quantization points. This gives equivalence of the distortion minimization problem to the Wasserstein minimization problem, when restricted to measures of finite support (Pagès, 2015).
@@
Suppose we have a quantization ${\mathbf{x}}=\{x_{1},...,x_{K}\}$. Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi cells $\mu(\partial C_{i})=0$. Then the measure $\nu$ that minimizes the Wasserstein-2 distance 14 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$, where $\delta(x_{i})$ denotes the Dirac delta distribution at $x_{i}$. Moreover, the optimal coupling is given by the projection onto the centroids.
Proof.
Deferred to Section˜C.2. ∎
In other words, finding points that minimize the quadratic distortion is equivalent to finding a $K$-finitely supported (probability) measure, minimizing the Wasserstein-2 distance to the underlying measure.
@@
The case where the approximating measure is a uniform Dirac mixture is called the Wasserstein barycenter problem (Cuturi and Doucet, 2014). The Wasserstein barycenter has higher error than the optimal quantization, but it admits an easily computable dual representation.
The quantizer can be shown to have nice approximation properties when taking expectations of functions. A prototypical example for DD would have $f$ be the gradient of a neural network with respect to some loss function. This implies that a data distribution $\mu$ and its quantization $\nu$ induce similar training dynamics.
@@ 2.1.1 Solving the optimal quantization problem
To find an optimal quantizer, we use the competitive learning vector quantization (CLVQ) algorithm (Ahalt et al., 1990). This arises directly from gradient descent on the quadratic distortion. The gradient of the distortion has a representation in terms of $\mu$-centroids of the corresponding Voronoi cells, given explicitly in Section˜C.5.
The CLVQ algorithm is presented in the appendix as Algorithm˜2. It consists of iterating: (i) sampling from $\mu$, (ii) computing the nearest cluster centroid, and (iii) updating the cluster centroid with weighted average. CLVQ is equivalent to the mini-batch $k$-means method when the step-sizes $\gamma_{i}$ are chosen to be the reciprocals of the number of points per-cluster (Sculley, 2010; Pedregosa et al., 2011). We can thus interpret mini-batch $k$-means as finding a local minima of the optimal quantization problem.
The CLVQ algorithm produces points ${\mathbf{x}}=(x_{1},...,x_{K})$, but it remains to compute the associated weights approximating the measures of the Voronoi cells $\mu(C_{i})$ as in Proposition˜1. This can be done in an online manner within the same iterations (Pagès, 2015).
Proposition 3 (Bally and Pagès 2003, Prop. 7).
Assume that the measure $\mu\in{\mathcal{P}}_{2+\eta}(\mathbb{R}^{d})$ for some $\eta>0$, and that it assigns no mass to hyperplanes. Assume further that the grids ${\mathbf{x}}^{(t)}$ produced by CLVQ converge to a stationary grid $\mathbf{x}$$^{*}$, i.e. $\nabla{\mathcal{G}}({\mathbf{x}}^{*})=0$, and that the step-sizes satisfy $\sum\gamma_{k}=+\infty$ and $\sum\gamma_{k}^{1+\delta}<\infty$ for some $\delta>0$. Then,
@@
2.2 Score-based diffusion
To connect the quantization error on the latent space with the quantization error in the image space, we need to consider properties of the latent-to-image process. In particular, we focus on score-based diffusion models, seen as discretizations of particular noising SDEs (Song et al., 2020). Consider the following SDE, where $W$ is a standard Wiener process:
[…]
The reverse of the diffusion process is given by the reverse-time SDE (Haussmann and Pardoux, 1986), where the density of $x$ at time $t>0$ is given by $p_{t}$,
[…]
and $\bar{W}$ is a reverse time Wiener process. For an increasing noising schedule $\sigma(t)$ or noise-scale $\beta(t)$, the variance-exploding SDE (VESDE, or Brownian motion) and variance-preserving SDE (VPSDE, or Ornstein–Uhlenbeck process) are given respectively by
[…]
These SDEs are related to early diffusion models. Specifically, VPSDE corresponds to denoising score matching with Langevin dynamics (Song and Ermon, 2019), and VESDE to denoising diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020). Using this particular structure, we may obtain convergence results as seen in the next section. We note that by time-rescaling, we may assume without loss of generality that the noising schedule is linear $\sigma^{2}(t)=t$ or the noise-scale is constant $\beta(t)=1$.
The goal in question: given Wasserstein-2 convergence of the marginals $\nu_{T}^{(k)}\rightarrow\mu_{T}$, we wish to derive a bound on expectations $\mathbb{E}_{\nu_{\delta}^{(k)}}[f]\xrightarrow{?}\mathbb{E}_{\mu_{\delta}}[f]$, for some small fixed $\delta\in(0,T)$ and $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying some regularity conditions. In other words, we wish to show that generative diffusion preserves closeness of data distributions. Such a bound would directly link to training neural networks with surrogate data, e.g. by taking $f$ to be the gradient of a loss function with respect to some network parameters.